# Market Making without Regret

### v1

@@ Appendix A Notation
Lemma C.2.
@@
The minimum value of $f$ is found when approaching $\nicefrac{{3}}{{16}}$ from the right, which yields $f(c)\rightarrow\nicefrac{{1}}{{6}}$, therefore $F(a)-F(b)\geq\nicefrac{{1}}{{6}}\cdot(a-b)$. ∎
Lemma C.3.
@@
(($\star$)$\leq$( ⋆ ) ≤ $\log\left(\frac{F(a)-F(b)}{F(a)-F(b)-\frac{\varepsilon_{K}}{18}\Lambda(b)}\right)(F(a)-F(b))+\log\left(\frac{F(b)}{F(b)+\frac{\varepsilon_{K}}{18}\Lambda(b)}\right)F(b)$
@@
    which is maximized when $\frac{\Lambda(a)-\Lambda(b)}{F(a)-F(b)}$ is minimized, by leveraging the Lipschitzness of $\Lambda$ from Lemma C.1 together with the lower bound on $F(a)-F(b)$ from Lemma C.2 we get
@@
  • Next consider the opposite $a-b>\varepsilon_{K}$. By Lemma C.2 we know that $F(a)-F(b)\geq\nicefrac{{1}}{{6}}(a-b)\geq\nicefrac{{\varepsilon_{K}}}{{6}}$. We use the inequality $\log x\leq x-1$ to get:
@@
In conclusion, the result holds with $c_{1}=\nicefrac{{2}}{{81}}$ and $c_{2}=\nicefrac{{1}}{{9}}+\nicefrac{{64}}{{9}}=\nicefrac{{65}}{{9}}$. ∎
Appendix D DKW inequalities for Section 4.3
In this section we present two bivariate DKW inequalities that can be deduced as corollaries of the VC-dimension theory [1, Theorem 4.9; see also Lemmas 4.4, 4.5, and 4.11 for the explicit constants].
Theorem D.1.
There exist positive constants $m_{0}\leq 1200$, $c_{1}\leq 13448$, $c_{2}\geq\nicefrac{{1}}{{576}}$ such that, if $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space, $(X_{n},Y_{n})_{n\in\mathbb{N}}$ is a $\mathbb{P}$-i.i.d. sequence of two-dimensional random vectors, then, for any $\varepsilon>0$ and all $m\in\mathbb{N}$ such that $m\geq\nicefrac{{m_{0}}}{{\varepsilon^{2}}}$, it holds
$\mathbb{P}\left[\sup_{x,y\in\mathbb{R}}\left\lvert\frac{1}{m}\sum_{k=1}^{m}\mathbb{I}\{X_{k}\leq x,Y_{k}\leq y\}-\mathbb{P}\left[X_{1}\leq x,Y_{1}\leq y\right]\right\rvert>\varepsilon\right]\leq c_{1}e^{-c_{2}m\varepsilon^{2}}\;,$
and
$\mathbb{P}\left[\sup_{x,y\in\mathbb{R}}\left\lvert\frac{1}{m}\sum_{k=1}^{m}\mathbb{I}\{X_{k}\leq x,Y_{k}<y\}-\mathbb{P}\left[X_{1}\leq x,Y_{1}<y\right]\right\rvert>\varepsilon\right]\leq c_{1}e^{-c_{2}m\varepsilon^{2}}\;.$

### v2

@@ Appendix B Additional Related Works
The following lemma, together with Lemma D.2, will be used in the proof of Lemma D.6.
Lemma D.4.
@@
The minimum value of $f$ is found when approaching $\nicefrac{{3}}{{16}}$ from the right, which yields $f(c)\rightarrow\nicefrac{{1}}{{6}}$, therefore $F(a)-F(b)\geq\nicefrac{{1}}{{6}}\cdot(a-b)$.
We conclude this section by proving a key lemma used in the proof of Theorem 4.1.
Lemma D.6.
@@
$(\star)\leq\log\left(\frac{F(a)-F(b)}{F(a)-F(b)-\frac{\varepsilon_{K}}{18}\Lambda(b)}\right)(F(a)-F(b))+\log\left(\frac{F(b)}{F(b)+\frac{\varepsilon_{K}}{18}\Lambda(b)}\right)F(b)$
@@
    which is maximized when $\frac{\Lambda(a)-\Lambda(b)}{F(a)-F(b)}$ is minimized, by leveraging the Lipschitzness of $\Lambda$ from Lemma D.2 together with the lower bound on $F(a)-F(b)$ from Lemma D.4 we get
@@
  • Next consider the opposite $a-b>\varepsilon_{K}$. By Lemma D.4 we know that $F(a)-F(b)\geq\nicefrac{{1}}{{6}}(a-b)\geq\nicefrac{{\varepsilon_{K}}}{{6}}$. We use the inequality $\log x\leq x-1$ to get:
@@
In conclusion, the result holds with $c_{1}=\nicefrac{{2}}{{81}}$ and $c_{2}=\nicefrac{{1}}{{9}}+\nicefrac{{64}}{{9}}=\nicefrac{{65}}{{9}}$.
Appendix E Proof of Theorem 5.1
  • $\bullet$
  • $\bullet$
  • $\bullet$
$R_{T}\geq\mathbb{E}\left[\sum_{t=1}^{T}u(b,a,M_{t},V_{t})\right]-\mathbb{E}\left[\sum_{t=1}^{T}u(B_{t},A_{t},M_{t},V_{t})\right]\\ \geq\left(\frac{1}{2}-\varepsilon\right)\cdot T-\left(\left(\frac{1}{2}+\varepsilon\right)\cdot\frac{1}{2}-\left(\frac{1}{2}-\varepsilon\right)\cdot\frac{1}{2}\right)\cdot T=\left(\frac{1}{2}-2\varepsilon\right)\cdot T\;.$
Setting $\delta\coloneqq 2\varepsilon$ and noticing that $\varepsilon$ was chosen arbitrarily in the interval $(0,\nicefrac{{1}}{{4}})$, the conclusion follows.